OpenStudy (anonymous):

write the equation of the line that goes through (3,5) and has a slope of 1/3

OpenStudy (anonymous):

use this : y-y1=m(x-x1), m=slope and (x1,y1)=(3,5)

OpenStudy (anonymous):

y= 1/3x + 0

OpenStudy (anonymous):

y= 1/3x+ 4 right?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

y= 1/3x+ 4 right?

OpenStudy (anonymous):

yes cause it would be y-5 = 1/3(x-3)

OpenStudy (anonymous):

omfg oops

OpenStudy (anonymous):

ur right lol

OpenStudy (anonymous):

haha ok maybe next time pal